OpenStudy (anonymous):

(2b)^-5 use positive exponents

jimthompson5910 (jim_thompson5910):

Similar Problem \[\Large (3y)^{-4} = \frac{1}{(3y)^{4}}\] \[\Large (3y)^{-4} = \frac{1}{(3^4)(y^{4})}\] \[\Large (3y)^{-4} = \frac{1}{81y^{4}}\]